# Image as an IMU: Estimating Camera Motion from a Single Motion-Blurred Image

### v3

@@ 5.2 Real-world Dataset
As our method, including the velocity solver, is fully-differentiable we can train it in an end-to-end manner. This enables us to additionally train on real-world motion blur to further close the reality-gap between our synthetically blurred dataset and in-the-wild images. We therefore collect 10k real-world motion-blurred images with corresponding ARKit poses, IMU measurements, and exposure times across about 30 scenes. We follow Liu et al. [28] and perform the Fast Fourier Transform to identify blurred images, where images within a threshold are blurry enough to be used. We also supplement with a small amount of images without blur to enhance the robustness to still frames.
@@ 6 Experiments
Our model is trained in three stages. We first train the flow and depth decoders without pose supervision on our synthetic dataset with a batch size of 32. Pose supervision is added once the depths and flow estimates are reasonable, and we train poses, flow and depths with a batch size of 8 for 300k steps. Finally, we finetune the model on the real-world collected images to close the reality gap from synthetic images for 10k steps. We train our method using Adam [20] on one Nvidia RTX 3090.
@@
To assess the accuracy of our method, we evaluate the rotational velocities and translational velocities in four unseen, real-world videos (see Supplementary 2). For each video sequence, we get the ARKit-estimated camera poses, IMU measurements, and the exposure times for each image. We directly compare predictions to the gyroscope measurements and predictions to a finite-difference velocity approximation from the ARKit translations. We treat the ARKit translations as the ground-truth as their visual-inertial method fuses additional sensor information, which helps mitigate the poor visual input. We measure the RMSE across each axis for and .
@@
There are no existing methods that can estimate camera motion from a single blurred image, so we instead use camera pose estimation methods and estimate the relative pose (and hence velocity) between frames using a centered finite-difference approximation. We evaluate against the following SfM and SLAM baselines:
COLMAP [39] [40] is an incremental-SfM method that is considered to be the de-facto method for obtaining camera poses and commonly serves as ground-truth for NeRFs [30] and Gaussian Splatting [19] methods. Because standard SfM is unreliable with small triangulation angles, we run COLMAP using all images in the sequence. After obtaining the camera poses across all images, we scale the trajectory with the ground-truth ARKit poses. We use the hloc package [38] to run COLMAP with SIFT [29] and AdaLAM [6] as well as a learning baseline using DISK [49] and LightGlue [26].
MASt3R [54] [23] is a recent learning-based 3D vision method that regresses the pointmaps between a pair of images. MASt3R is the current SOTA in retrieving metric-scale camera poses, particularly with extremely challenging views with little-to-no-overlap. To better represent our method, we evaluate MASt3R on pairwise images two frames apart.
DROID-SLAM [47] is a learning-based SLAM method that utilizes RAFT [46] to estimate the optical flow across frames to help compute the camera poses, and obtains SOTA accuracy and robustness particularly in adversarial conditions. To obtain the camera poses for each frame, we set DROID-SLAM to identify each image as a keyframe by setting the filter and keyframe thresholds to 0.
Note that all baselines require two or more images to estimate the velocity, whereas our method only needs a single image. We therefore tackle a more difficult problem, and our formulation also more closely resembles the true camera state.
| Method | Runtime (s) | Realtime factor |
| COLMAP (D+LG) | 182.98 | 0.038 |
| MASt3R | 67.90 | 0.103 |
| DROID-SLAM | 41.61 | 0.169 |
| Ours | 6.48 | 1.08 |
@@ Results.
We report the rotational and translational RMSE in Tables 1 and 2. While the recorded video sequences suffer from severe motion blur, our method is able to handle the motion very robustly. Figure 5 highlight the overall robustness of our method, with the error CDFs demonstrating that our method suffers from noticably less errors than the baselines. In these sequences, the baseline methods suffer from major outliers and zero-prediction failures, which can be observed in Figure 4. Velocity predictions from the billiards sequence are shown with the remaining sequences included in Supplementary 3. Overall, we obtain an average of 31% reduction of RMSE from the second-best result in rotational velocities and a 24% reduction in translational velocities.
To further showcase the usefulness of our method, we test our method in estimating the velocity of the gripper on a moving robot arm. We place the camera on the robot pictured in Figure 6(a) and predict the velocity of the arm solely from the images in the recorded video. The motion-blurred images give a salient enough signal to accurately estimate the movement of the arm as seen in Figure 6(b).
In Figure 7 we show a comparison between our method and an actual IMU. To do this test we first estimate the IMU bias from a stationary calibration video, followed by estimating the initial pose of the camera to properly align with the gravity vector. We then preintegrate the IMU measurements following [13] using GTSAM [11] and compare the integrated velocities to our predictions. Figure 7 shows the drift that occurs in and after only 20 seconds, whereas our method is reliably drift-free for the entire video. Therefore, while our method produces inertial-like measurements, we see in this example that our method provides an even better motion estimate than an IMU.
We lastly measure the runtime of our method. Even when running on every frame in the sequence, we can run in real-time at 30 FPS, where the inference and direction disambiguation for each image takes on average 0.03 seconds on an RTX 3090 GPU.

### v4

@@ 5.2 Real-world Dataset
As our method, including the velocity solver, is fully-differentiable, we can train it in an end-to-end manner. This enables us to additionally train on real-world motion blur to further close the reality gap between our synthetic dataset and in-the-wild images. We collect 10k real-world motion-blurred images with corresponding ARKit poses, IMU measurements, and exposure times across about 30 scenes. We follow Liu et al. [liu2008cvpr_blur-detect] and perform the Fast Fourier Transform to identify blurred images, where images within a threshold are blurry enough to be used. We also add images without blur to enhance the robustness to still frames.
@@ 6 Experiments
Our model is trained in three stages. We first train the flow and depth decoders without pose supervision on our synthetic dataset with a batch size of 32. Pose supervision is added once the depths and flow estimates are reasonable, and we train poses, flow and depths with a batch size of 8 for 300k steps. Finally, we finetune the model on real-world collected images for 10k steps. We train our method using Adam [kingma2017adam] on an Nvidia RTX 3090 GPU.
@@
To assess the accuracy, we evaluate the rotational velocities and translational velocities in four unseen, real-world videos. These test sequences explicitly have different domains from those in training and have varying exposure times to test its generalization. We do not evaluate on existing motion-blurred benchmarks such as GoPro [Nah_2017_CVPR] since neither poses nor velocities are provided, and computing poses using COLMAP and MASt3R were inaccurate due to small baselines. For each test sequence, we get the ARKit-estimated camera poses, IMU measurements, and the exposure times for each image. We directly compare predictions to the gyroscope measurements and predictions to a finite-difference velocity approximation from the ARKit translations. We use these translations as our ground-truth since ARKit’s visual-inertial method tightly fuses IMU measurements and is robust even with poor visual tracks. We also only use the computed ARKit instantaneous body-frame velocity, mitigating any major drift. We do not use COLMAP as we found it unreliable in these scenarios (see Tab. 1 and 2). We measure the RMSE across each axis for and .
@@
No existing methods estimate camera motion from a single blurred image, so we instead use SfM/SLAM methods and estimate the relative pose (and hence velocity) between frames using a centered finite-difference approximation. We evaluate against the following baselines:
COLMAP [schoenberger2016sfm] [schoenberger2016mvs] is an incremental-SfM method that is considered the de-facto method for obtaining camera poses. Because standard SfM is unreliable with small triangulation angles, we run COLMAP with all sequence images. After estimating the trajectory, we scale it using the ground-truth ARKit poses. We use hloc [sarlin2019coarse] to run COLMAP with SIFT [Lowe2004DistinctiveIF] + AdaLAM [cavalli2020handcrafted] as well as a learned baseline with DISK [tyszkiewicz2020disk] + LightGlue [lindenberger2023lightglue].
MASt3R [dust3r_cvpr24] [mast3r_arxiv24] is a recent learning-based 3D vision method that regresses the pointmaps between a pair of images. MASt3R is the current SOTA in retrieving metric-scale camera poses, particularly with sparse views with little-to-no overlap. To better compare to our method, we evaluate MASt3R on pairwise images two frames apart.
DROID-SLAM [teed2021droid] is a learning-based SLAM method that uses RAFT [teed_raft_2020] to estimate the optical flow across frames to help compute the camera poses. To obtain the camera poses for each frame, we set the filter and keyframe thresholds to 0.
We also provide the zero-velocity baseline representing a stationary camera, which serves as a reasonable prior for most trajectories. Note that all baselines require two or more images to estimate the velocity, whereas our model only ever sees a single image. We therefore tackle a more difficult problem, and our formulation more closely resembles the true camera state.
| Method | Average FPS |
| COLMAP (D+LG) | 0.95 |
| MASt3R | 2.83 |
| DROID-SLAM | 4.34 |
| Ours | 30.17 |
@@ Results.
We report the rotational and translational RMSE in Tables 1 and 2. While the recorded video sequences suffer from severe motion blur (see supplementary for examples), our method is able to handle the motion robustly. Figure 5 highlights the overall robustness of our method, with the error CDFs demonstrating that our method suffers from noticably less errors than the baselines. In these sequences, the baseline methods suffer from major outliers and zero-prediction failures, observed in Figure 4. Overall, we obtain an average of 31% reduction of RMSE from the second-best result in rotational velocities and a 24% reduction in translational velocities.
We also test our method in a robotics application by estimating the gripper velocity on a robot arm. We place the camera on the robot pictured in Figure 6(a) and predict the gripper velocity solely from its video. The motion-blurred images give a salient enough signal to accurately estimate the movement of the arm as seen in Figure 6(b).
In Figure 7 we compare our method with using an actual IMU. We first estimate the IMU bias from a stationary calibration video, followed by estimating the initial pose of the camera to properly align with the gravity vector. We then preintegrate the IMU measurements following [forster2016manifold] using GTSAM [gtsam] and compare the integrated velocities to our predictions. Figure 7 shows the drift that occurs in and after only 20 seconds, whereas our method is reliably drift-free for the entire video. While we are comparing our direct translational velocity estimates to the IMU’s preintegrated acceleration, we argue that this is advantageous for state estimation since we bypass the drift-accruing integration. Therefore, we can see how our method can provide an even better motion estimate than an IMU.
We lastly measure each method’s average FPS (Table 3). Even when running on every frame, our method can run in real-time at 30 FPS, with inference and direction disambiguation taking 0.03 seconds on an RTX 3090 GPU.